OpenStudy (anonymous):

Can someone please help me with this equation. Find the derivative of the function. h(x) = 5/x^2 + 5/x^3 h'(x)= ??

zepdrix (zepdrix):

Just rewrite h(x) using negative exponents :O should make it alil easier to see what's going on. Just the power rule from there. \[\huge h(x)=5x^{-2}+5x^{-3}\]

OpenStudy (anonymous):

-10x-15x^2?

zepdrix (zepdrix):

Remember that when you SUBTRACT in the negative direction, then numbers should actually get larger. You're subtracting one from the power, here is what the first term will look like. \[\huge (5x^{-2})'=5(-2)x^{-2-1}=-10x^{-3}\]

OpenStudy (anonymous):

ok, then the second term would be -15x^-4

zepdrix (zepdrix):

mhm c:

OpenStudy (anonymous):

so -10x^-3-15x^-4?

zepdrix (zepdrix):

yes, gj

OpenStudy (anonymous):

@zepdrix can you help me with one more problem?

zepdrix (zepdrix):

sure

OpenStudy (anonymous):

\[ = (x^2-3x-1)^-7\]

zepdrix (zepdrix):

So you'll take the derivative of the outermost function first, which is ( )^-7

zepdrix (zepdrix):

Then we have to apply the chain rule :O multiply by the derivative of the inner function.

zepdrix (zepdrix):

\[\large \frac{ d }{ dx }(x^2-3x-1)^{-7}=-7(x^2-3x-1)^{-8} [\frac{ d }{ dx }(x^2-3x-1)]\]

zepdrix (zepdrix):

Confused? :O maybe the prime notation is a little easier to read, i dunno :3

OpenStudy (anonymous):

just going over what you said.

OpenStudy (anonymous):

ok i got it, so what would the next steo be?

zepdrix (zepdrix):

\[\large \frac{ d }{ dx }(x^2-3x-1)\] Hmm looks like we need to take this derivative still.

OpenStudy (anonymous):

2x-3

zepdrix (zepdrix):

Mmmmmmmmmmmmmm yah sounds good!

zepdrix (zepdrix):

\[\huge -7(x^2-3x-1)^{-8}(2x-3)\] and yah that should be our final answer, yay.

OpenStudy (anonymous):

thanks a lot zep

OpenStudy (anonymous):

one more? lol